OpenStudy (anonymous):

Which of the following is greatest? seven twenty-fifths zero point three three nineteen one hundredths eleven-fiftieths

OpenStudy (turingtest):

can you long-divide "seven twenty-fifths" ? 7/25=?

OpenStudy (turingtest):

do that with all, then compare

OpenStudy (turingtest):

the third choice is already in decimal form

OpenStudy (anonymous):

well i still dnt understand

OpenStudy (turingtest):

do you not know how to do long division?

OpenStudy (turingtest):

there are other ways to do it, but I think that is the most direct @Nick_Black please don't post direct answers without explain, even if you write "maybe"

OpenStudy (anonymous):

You should make them all into fractions with common denominators and then try.

OpenStudy (anonymous):

Mr TT i wrote direct bcoz it very easy and its understanible from question tht u have to fraction them all and the greatest one will be know for this i also need to explain????

OpenStudy (anonymous):

@Nick_Black it states in CoC that we are not allowed to just give an answer as that promotes cheating. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

okay my bad i will explain it 7/25=0.28 19/100=0.19 11/50=0.22 therefore 0.33 is the greatest of all

OpenStudy (anonymous):

rebecca if explaination was needed viv would have asked for it i would not deny it

OpenStudy (anonymous):

and how the hell u tag me????

OpenStudy (turingtest):

I know it is easier to just give the answer. that is the challenge: to teach instead

OpenStudy (turingtest):

and please don't curse

OpenStudy (anonymous):

She shouldn't need to ask for explanation. You should be giving it or not replying. and I just put "@" at the beginning of your username.

OpenStudy (karatechopper):

Vivi i dont understand what other help u needed in this problem. There were 3 people who helped ya out up there^

OpenStudy (karatechopper):

Sorry 2

OpenStudy (anonymous):

She's offline

OpenStudy (karatechopper):

She was looking around just a few mins ago!!!!

OpenStudy (anonymous):

i know i just need explanation

OpenStudy (karatechopper):

Look at nick blacks